OpenStudy (ajanijones):

Determine the sum of the given probabilities. Are the events complementary? P(red) =2/5 ; P(blue) = 3/5 A. 1/2 The events are complementary. B. 1/2 The events are not complementary. C. 1 The events are complementary. D. 1 The events are not complementary.

OpenStudy (ajanijones):

PLease help if you can please.

OpenStudy (ajanijones):

be right back

OpenStudy (anonymous):

P(red) +P(blue) =2/5+ 3/5 =1

OpenStudy (ajanijones):

Ok so is in C or D.

OpenStudy (anonymous):

The events are complementary.

OpenStudy (anonymous):

because P(red) +P(blue) =1